OpenStudy (anonymous):

Find all the complex roots. Write the answer in the indicated form. The complex cube roots of 27(cos 234° + i sin 234°) (polar form) A. -3(cos 78° + i sin 78°), 3(cos 198° + i sin 198°), -3(cos 318° + i sin 318°) B. 3(cos 78° + i sin 78°), 3(cos 118° + i sin 118°), 3(cos 158° + i sin 158°) C. -3(cos 78° + i sin 78°), 3(cos 118° + i sin 118°), -3(cos 158° + i sin 158°) D. 3(cos 78° + i sin 78°), 3(cos198° + i sin 198°), 3(cos 318° + i sin 318°) I know it is not A I got this one wrong

OpenStudy (anonymous):

i think the answer is D

OpenStudy (anonymous):

the cube root of 27 is 3 and \(243\div3=78\)

OpenStudy (anonymous):

go around the circle again \((234+360)\div 3=198\) and again \((234+720)\div 3=318\)

OpenStudy (anonymous):

looks like D

OpenStudy (anonymous):

thank you again i always second guess myself

OpenStudy (anonymous):

when you write a number in polar form as \(r\left(\cos(\theta)+i\sin(\theta)\right)\) \(r=\sqrt{a^2+b^2}\) should be positive

OpenStudy (anonymous):

yw